a clock reads 5:30
a) what is the radian measure of the smaller angle between the hour and the minute hand
b) what is the linear speed of the tip of the minute hand on a clock if that hand is 5cm long

Answer by TimothyLamb(4379)

a clock reads 5:30
---
what is the radian measure of the smaller angle (A) between the hour and the minute hand?
---
the hour hand is at h radians clockwise from 12 noon
h = (5/12) * 2 * pi ~= 2.618
---
the minute hand is at m radians clockwise from 12 noon
m = (6/12) * 2 * pi ~= 3.142
---
A = 3.142 - 2.618 = 0.524 radians
---
what is the linear speed of the tip of the minute hand on a clock if that hand is 5 cm long
---
s = d / t
---
C = 2 * pi * R = 2 * pi * 0.05 = 0.31416 m/rev
---
angular speed = (1 rev / 60 min) * (1 min / 60 sec) = 1/3600 rev/sec
---
linear tip speed = (0.31416 m/rev) * (1/3600 rev/sec) = 8.727 * 10^-5 m/sec
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php